OMX service has been performed, GUI bug workaround

The OMX has been serviced and new OTFs have been made. We recommend using these. A strange bug in the acquisition software causes the 405 nm laser line (1st channel) to present wrong values, i.e. intensity % of the laser in the drop-down list do not represent the actual values. The desktop background picture of the PC instructs you restore the right interface (in 3 clicks).